OpenStudy (anonymous):

x^2-7x+5=0 Find the roots of the quadractic equation and obtain their sum and product.

OpenStudy (anonymous):

\[x=\frac{-(-7) \pm \sqrt{(-7)^2-4(1)(5)}}{2(1)}\]

OpenStudy (anonymous):

we can do the second question before the first

OpenStudy (anonymous):

the product is 5, and the sum is 7 you can read that right off the the equation

OpenStudy (anonymous):

so what if it a variable like ax^2+bx+c=0

OpenStudy (anonymous):

\[x=\frac{7 \pm \sqrt{49-20}}{2}=\frac{7 \pm \sqrt{29}}{2}\]

OpenStudy (anonymous):

\[\frac{7+\sqrt{29}}{2}+\frac{7-\sqrt{29}}{2}=\frac{14}{2}=7\]

OpenStudy (anonymous):

@monkey131 i dont think they are looking for that

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

they jusy wana fine the a+B and aB

OpenStudy (anonymous):

you knew it was 7 in advance. also you know the product is 5 if you have \[x^2+bx+c=0\] and the roots are \[r_1,r_2\] then \[r_1\times r_2=c\] and \[r_1+r_2=-b\]

OpenStudy (anonymous):

\[\frac{7-\sqrt{29}}{2}\cdot \frac{7+\sqrt{29}}{2}=\frac{49-29}{2}=\frac{20}{10}=2\]

OpenStudy (anonymous):

a + b = -b/a ab = c/a

OpenStudy (anonymous):

how'd u make it 49

OpenStudy (anonymous):

@monkey

OpenStudy (anonymous):

7 times 7

OpenStudy (anonymous):

@Satellite : im still dont get it T.T

OpenStudy (anonymous):

@monky denominator should be 4 product is 5

OpenStudy (anonymous):

Which part of the question don't you get?

OpenStudy (anonymous):

you are right satellite so we have 20/4=5

OpenStudy (anonymous):

@killer, if you have \[x^2+bx+c=0\] and the roots are \[r_1,r_2\] then it factors as \[(x-r_1)(x-r_2)\]

OpenStudy (anonymous):

If you complete the square of ax^2 + bx + c = 0 it would explain it better..I think

OpenStudy (anonymous):

when you multiply out you get \[x^2+bx+c=x^2-r_1x-r_2+r_1\times r_2\] \[=x^2-(r_1+r_2)x+r_1r_2\]

OpenStudy (anonymous):

therefore \[r_1r_2=c\] and \[r_1+r_2=-b\]

OpenStudy (anonymous):

ohh i see isee thank you :DD FINALY LOL

OpenStudy (anonymous):

and nayedo is right, you will get the same thing if you use the quadratic formula and multiply out

OpenStudy (anonymous):

shouldnt the sum be '-7'?

OpenStudy (anonymous):

so lets use the quadratic equaltion i got to the part where 7+- (square root of 49-20)/2

OpenStudy (anonymous):

a + b = \[(7 + \sqrt{29})/ 2 + (7-\sqrt{29})/ 2 = 14 / 2 = 7\]

OpenStudy (anonymous):

\[ab = (7+\sqrt{29})/2 \times (7-\sqrt{29})/2 = (49 - 29)/4 = 5\]

OpenStudy (anonymous):

@jahtoday it is \[x^2+bx+c=x^2-(r_1+r_2)x+r_1r_2\]so \[r_1+r_2=-b\]

OpenStudy (anonymous):

yup u r correct